# Norming Markushevich bases: recent results and open problems

### v1

@@ 1. Introduction
where the series is assumed to converge in the norm of $\mathcal{X}$. This permits an identification between the vector $x\in\mathcal{X}$ and the scalar sequence $(x_{k})_{k=1}^{\infty}$ and allows to view the Banach space $\mathcal{X}$ as a more concrete object whose elements are sequences of reals. Among the advantages of this representation, let us mention for instance that algebraic manipulations become more handy and linear operators can be represented by infinite matrices. Accordingly, Schauder bases have been extensively exploited in Banach space theory to obtain several structure results, see, e.g., [1, 31, 88]. Despite their usefulness, they have at least two drawbacks: first, the existence of a Schauder basis in $\mathcal{X}$ clearly implies that $\mathcal{X}$ must be separable; second, even though all classical separable Banach spaces admit one, there are separable Banach spaces without Schauder bases, [24].
@@
A well-known, yet non-trivial, result due to Banach is that the coefficient functionals $\varphi_{k}\colon\mathcal{X}\to\mathbb{R}$ given by $x\mapsto x_{k}$ are continuous. Additionally, ${\rm span}\{e_{k}\}_{k=1}^{\infty}$ is dense in $\mathcal{X}$ and the functionals $\{\varphi_{k}\}_{k=1}^{\infty}$ separate points on $\mathcal{X}$. Therefore, one can consider the following generalisation of a Schauder basis. A sequence $\{e_{k};\varphi_{k}\}_{k=1}^{\infty}\subseteq\mathcal{X}\times\mathcal{X}^{*}$ is a Markushevich basis (hereinafter M-basis) if $\langle\varphi_{k},e_{n}\rangle=\delta_{k,n}$, ${\rm span}\{e_{k}\}_{k=1}^{\infty}$ is dense in $\mathcal{X}$, and ${\rm span}\{\varphi_{k}\}_{k=1}^{\infty}$ is $w^{*}$-dense in $\mathcal{X}^{*}$. This more general definition solves at once the two drawbacks of Schauder bases from the previous paragraph: indeed, it is clear how to formulate the definition of M-basis in a way that would not require separability and M-bases are known to exist in all separable Banach spaces [75] (we will say more on this in Section 2). In addition, there are natural examples of M-bases that are not Schauder bases, such as, for example, the classical trigonometric system from Fourier Analysis, see Example 2.2. More importantly, M-bases have been successfully exploited to prove results valid for all separable Banach spaces; as a small sample, let us mention [6, 21, 44, 46, 50]. Before we continue, we should mention the main weakness of M-bases compared to Schauder bases: while it remains true that every vector $x\in\mathcal{X}$ can be identified with its coordinates $(\langle\varphi_{k},x\rangle)_{k=1}^{\infty}$, the series $\sum_{k=1}^{\infty}\langle\varphi_{k},x\rangle e_{k}$ does not converge in general.
@@
However, it is the non-separable framework the one where M-bases have proved being the most fruitful. As a first instance of this, let us mention that every Banach space with an M-basis admits a continuous injection in $c_{0}(\Gamma)$ and therefore a rotund renorming, see, e.g., [23, Theorem II.2.4]. In order to achieve stronger renorming properties one frequently requires additional assumptions on the M-basis. Perhaps the most important example is the celebrated Troyanski’s renorming theorem [96] that every Banach space with a strong M-basis admits a locally uniformly rotund (LUR, for short) norm. Let us also refer to [43, Theorem 3.48] for a proof and to [19] and [77, Lemma 2.1] for related results. Among all properties of M-bases the most close one to renorming theory is that of norming M-basis, since the very definition of norming M-basis involves an equivalent norm induced by the M-basis itself (see Section 2 for details).
@@
A related use of M-bases in non-separable Banach spaces is to characterise several classes of non-separable Banach spaces by the existence of M-bases with certain additional properties. The most important examples of this are weakly Lindelöf determined (WLD, for short) and Plichko spaces that we will define in Section 3 and will play a role in several parts of the paper. Let us refer to [43, Chapter 6] and [103] for more on this perspective; let us also mention in passing a parallel line of research aiming at characterising classes of Banach spaces via certain properties of so called projectional skeletons, see, e.g., [15, 18, 32].
@@
One of the main open problems in the area, still widely open, has been to detect which class of Banach spaces is characterised by the existence of a norming M-basis. It has been conjectured in the Seventies that there ought to be a connection between norming M-bases and weakly compactly generated Banach spaces (WCG for short). This resulted in a famous open problem, due to John and Zizler [53], whether every WCG Banach space admits a norming M-basis and to a question in the opposite direction, due to Godefroy [39, Problem 112], whether Asplund spaces with norming M-bases are WCG. Both these problems have been recently solved in the negative [42, 45] and we shall give more details on this in Section 4.
@@
is easily seen to be a continuous injection (here, $[-1,1]^{\mathbb{N}}$ is endowed with the product topology). By compactness of $(B_{\mathcal{X}^{*}},w^{*})$, it is a homeomorphic embedding, which implies the standard fact that $(B_{\mathcal{X}^{*}},w^{*})$ is metrisable. The same evaluation map $\mathsf{ev}$, now having values in $[-1,1]^{\Gamma}$, can be defined in presence of any M-basis and it still is a homeomorphic embedding. Depending on the range of $\mathsf{ev}$ one can classify several classes of Banach spaces: for example, when $\mathcal{X}$ is WLD $\mathsf{ev}(\varphi)$ only has countably many non-zero coordinates, which leads to the notion of Corson compact space. Similar results are available when $\mathcal{X}$ is WCG, Hilbert generated, or Plichko, as we will explain in Section 3. When the M-basis is $1$-norming, the description of the image of $\mathsf{ev}$ leads to the notion of semi-Eberlein compact space, first introduced in [69]. In Section 7 we shall define this class of compact spaces, describe some known properties, and state several open problems.
@@ 2. Examples of M-bases
Given a Banach space $\mathcal{X}$, a family $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}\subseteq\mathcal{X}\times\mathcal{X}^{*}$ is a biorthogonal system in $\mathcal{X}$ if $\langle\varphi_{\alpha},e_{\beta}\rangle=\delta_{\alpha,\beta}$ for all $\alpha,\beta\in\Gamma$. The biorthogonal system $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is fundamental (or complete) if ${\rm span}\{e_{\alpha}\}_{\alpha\in\Gamma}$ is dense in $\mathcal{X}$ and it is total if ${\rm span}\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is $w^{*}$-dense in $\mathcal{X}^{*}$ (equivalently, if it separates points of $\mathcal{X}$: the unique vector $x\in\mathcal{X}$ with $\langle\varphi_{\alpha},x\rangle$ for all $\alpha\in\Gamma$ is $x=0$). A Markushevich basis (henceforth, M-basis) is a fundamental and total biorthogonal system. As we said in the Introduction, one frequently requires M-bases with additional properties, that are frequently obtained by strengthening the sense in which the functionals $\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ exhaust the dual. The strongest notion is that of shrinking M-basis, obtained by requiring that ${\rm span}\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is dense in $\mathcal{X}^{*}$ (in the norm topology); as we will see in Theorem 4.3 this is a rather restrictive definition.
@@
Observe that $\mathcal{Z}$ becomes a $1$-norming subspace in the norm $\left\|\cdot\right\|_{\mathcal{Z}}$. Let us refer to [40, Section 3.4] for some additional elementary remarks on the notion of norming subspaces.
@@
$\int_{\mathbb{T}}\alpha F_{n}(t_{0}-t)f(t)\ {\rm d}\mu(t)=\alpha(F_{n}*f)(t_{0})\to\|f\|_{\infty}$
@@
More generally, $\mathcal{C}(\mathcal{K})$ has a $1$-norming M-basis whenever $\mathcal{K}$ is a compact Abelian group, [64, Theorem 4.6].
@@
Every Schauder basis is a norming M-basis and every monotone Schauder basis is a $1$-norming one. This is elementary and we refer to [40, Remark 32(3)] for a proof. A similar argument shows that every long unconditional basis is a norming M-basis as well, [43, Exercise 7.8]. In particular, the canonical bases of $c_{0}(\Gamma)$ and $\ell_{p}(\Gamma)$, $1\leqslant p<\infty$ are $1$-norming. In the non-separable case the unconditionality assumption is essential, as not every long Schauder basis is norming (see Remark 5.4).
In the opposite direction, the classical Mazur technique shows that every norming M-basis admits a Schauder subsequence; we state this only for countable M-bases, but a small variation gives the result for uncountable ones as well, [45, Fact 4.4].
@@
and a standard inductive argument (as in [31, Theorem 4.19]) finishes the proof. ∎
Next, we discuss the existence of (norming) M-bases in separable Banach spaces. The following crucial result, essentially based on the Gram-Schmidt orthogonalisation algorithm, is due to Markushevich [75]. We refer to [31, Theorem 4.59] for a proof.
@@
[…]
The result immediately implies the existence of a $1$-norming M-basis in every separable Banach space. Indeed, it is enough to take a dense sequence $(x_{k})_{k=1}^{\infty}$, let $\psi_{k}$ be a norm one functional with $\langle\psi_{k},x_{k}\rangle=1$ and observe that ${\rm span}\{\psi_{k}\}_{k=1}^{\infty}$ is $1$-norming for $\mathcal{X}$.
The situation is radically different for non-separable $\mathcal{X}$ and there are non-separable Banach spaces without an M-basis, the simplest example of this being $\ell_{\infty}$, [58]. It is even consistent with ZFC that there are non-separable Banach spaces without uncountable biorthogonal systems, [43, Section 4.4].
Part of the interest of Theorem 2.5 is in the clause concerning the preservation of the linear span, that allows for instance to obtain examples of non-norming M-bases.
@@
When searching for M-bases that fail to be norming one immediately faces the problem of finding $w^{*}$-dense subspaces that are not norming. The existence of such subspaces has been clarified in [20]: A Banach space $\mathcal{X}$ has a $w^{*}$-dense, not norming, subspace in $\mathcal{X}^{*}$ if and only if $\mathcal{X}^{**}/\mathcal{X}$ is infinite-dimensional (i.e., $\mathcal{X}$ is not quasi-reflexive). Therefore, every separable Banach space that is not quasi-reflexive admits a non-norming M-basis. Conversely, in a separable quasi-reflexive Banach space all M-bases are norming (and in a separable reflexive space all M-bases are shrinking).
@@
For instance, the trigonometric system of Example 2.2 is strong. Despite how innocent this condition might look there are M-bases that fail to be strong and it is actually rather easy to find one in every separable Banach space, [43, Proposition 1.34]. Inspection of this proof shows that, if $\{e_{k};\varphi_{k}\}_{k=1}^{\infty}$ is an M-basis for $\mathcal{X}$, then there is another M-basis for $\mathcal{X}$ with the same linear spans and that fails to be strong. Starting from a norming M-basis shows that every separable Banach space admits a norming M-basis that is not strong. We will see an example of the opposite situation in Remark 5.6.
@@ 3. Classes of non-separable Banach spaces and M-bases
WCG Banach spaces admit a characterisation in terms of M-bases as follows: a Banach space $\mathcal{X}$ is WCG if and only if it admits a weakly compact M-basis, i.e., an M-basis $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ such that $\{e_{\alpha}\}_{\alpha\in\Gamma}\cup\{0\}$ is weakly compact in $\mathcal{X}$ (see [43, Theorem 6.9]). It is easy to reformulate this property in terms of the evaluation map $\mathsf{ev}$: the M-basis is weakly compact if and only if
@@
When $\mathcal{X}$ is a $\mathcal{C}(\mathcal{K})$ space, the situation is more symmetric: in fact $\mathcal{C}(\mathcal{K})$ is WCG if and only if $\mathcal{K}$ is Eberlein, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Eberlein, [5]. This and the fact that $\mathcal{X}$ is a subspace of $\mathcal{C}(B_{\mathcal{X}^{*}})$ imply that, when $B_{\mathcal{X}^{*}}$ is Eberlein, $\mathcal{X}$ is a subspace of the WCG space $\mathcal{C}(B_{\mathcal{X}^{*}})$. Therefore the dual ball of $\mathcal{X}$ is Eberlein if and only if $\mathcal{X}$ is a subspace of a WCG space.
At one point, we will need a subclass of the class of WCG spaces, namely Hilbert generated Banach spaces and the corresponding class of uniform Eberlein compacta. A Banach space $\mathcal{X}$ is Hilbert generated if there are a Hilbert space $\mathcal{H}$ and a linear map $T\colon\mathcal{H}\to\mathcal{X}$ with dense range. By definition, separable Banach spaces, $c_{0}(\Gamma)$ and $L_{1}(\mu)$ for a finite measure $\mu$ are Hilbert generated. However, there are reflexive Banach spaces that are not Hilbert generated [70] and even super-reflexive ones, [28, Section 4]. Since Hilbert spaces are WCG, every Hilbert generated Banach space is WCG. Dually, a compact topological space $\mathcal{K}$ is uniform Eberlein if it is homeomorphic to a weakly compact subset of a Hilbert space. Analogous results as those of the previous paragraphs give that $\mathcal{X}$ is Hilbert generated if and only if
@@
and that $\mathcal{X}$ is a subspace of a Hilbert generated space if and only if $(B_{\mathcal{X}^{*}},w^{*})$ is uniform Eberlein. Moreover, a $\mathcal{C}(\mathcal{K})$ space is Hilbert generated if and only if $\mathcal{K}$ is uniform Eberlein, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is uniform Eberlein, [12, Theorem 3.2]. For further details we refer to [43, Section 6.3] and [12, 28, 30].
@@
In this case, it turns out that every M-basis countably supports $\mathcal{X}^{*}$, [43, Theorem 5.37]. Comparing this notion to the M-basis characterisation of WCG spaces one easily sees that every WCG Banach space is WLD. Moreover, WLD spaces are closed under quotients and subspaces, as we will mention below (thus subspaces of WCG are also WLD). It is not easy to offer natural examples of WLD spaces that are not WCG; one could mention for instance Rosenthal’s non WCG subspace of a WCG space [83], or Talagrand’s example of a WLD space that is not subspace of a WCG space, [92].
@@
Therefore, $(B_{\mathcal{X}^{*}},w^{*})$ satisfies the following definition. A compact topological space $\mathcal{K}$ is Corson if it is homeomorphic to a subset of $\Sigma(\Gamma)$, for some set $\Gamma$. Since elements in $c_{0}(\Gamma)$ are countably supported, every Eberlein compact space is Corson. Moreover, Corson compacta are closed under subspaces and continuous images, [38, 41, 76]. As it turns out, a Banach space is WLD if and only if its dual ball is Corson, [7, Section 1], [100, Theorem 1.1] and references therein. Therefore, we get in particular that WLD spaces are closed under quotients and subspaces. Finally, a $\mathcal{C}(\mathcal{K})$ space is WLD if and only if $\mathcal{K}$ is Corson and it has property (M), namely every measure on $\mathcal{K}$ has separable support, [8].
At one point (Theorem 4.3) we will need one more class, intermediate between WCG and WLD Banach spaces. In order not to overload the section with definitions we prefer not to enter its definition and only give some references. The class consists of the so-called weakly countably determined (WCD, for short), or Vašák, Banach spaces; as we said, we have the implications WCG $\implies$ WCD $\implies$ WLD. WCD spaces were introduced in [101], see e.g., [23, Section VI.2], or [40, p. 174 and p. 433].

### v2

@@ 1. Introduction
where the series is assumed to converge in the norm of $\mathcal{X}$. This permits an identification between the vector $x\in\mathcal{X}$ and the scalar sequence $(x_{k})_{k=1}^{\infty}$ and allows to view the Banach space $\mathcal{X}$ as a more concrete object whose elements are sequences of reals. Among the advantages of this representation, let us mention for instance that algebraic manipulations become more handy and linear operators can be represented by infinite matrices. Accordingly, Schauder bases have been extensively exploited in Banach space theory to obtain several structural results, see, e.g., [1, 38, 99]. Despite their usefulness, they have at least two drawbacks: first, the existence of a Schauder basis in $\mathcal{X}$ clearly implies that $\mathcal{X}$ must be separable; second, even though every classical separable Banach space admits one, there are separable Banach spaces without Schauder bases, [30].
@@
A well-known, yet non-trivial, result due to Banach is that the coefficient functionals $\varphi_{k}\colon\mathcal{X}\to\mathbb{R}$ given by $x\mapsto x_{k}$ are continuous. Additionally, ${\rm span}\{e_{k}\}_{k=1}^{\infty}$ is dense in $\mathcal{X}$ and the functionals $\{\varphi_{k}\}_{k=1}^{\infty}$ separate points on $\mathcal{X}$. Therefore, one can consider the following generalisation of a Schauder basis. A sequence $\{e_{k};\varphi_{k}\}_{k=1}^{\infty}\subseteq\mathcal{X}\times\mathcal{X}^{*}$ is a Markushevich basis (hereinafter M-basis) if $\langle\varphi_{k},e_{n}\rangle=\delta_{k,n}$, ${\rm span}\{e_{k}\}_{k=1}^{\infty}$ is dense in $\mathcal{X}$, and ${\rm span}\{\varphi_{k}\}_{k=1}^{\infty}$ is $w^{*}$-dense in $\mathcal{X}^{*}$. This more general definition solves at once the two drawbacks of Schauder bases from the previous paragraph: indeed, it is clear how to formulate the definition of an M-basis in a way that would not require separability, and M-bases are known to exist in all separable Banach spaces [85] (we will say more on this in Section 2). In addition, there are natural examples of M-bases that are not Schauder bases, such as, for example, the classical trigonometric system from Fourier Analysis, see Example 2.2. More importantly, M-bases have been successfully exploited to prove results valid for all separable Banach spaces; as a small sample, let us mention [6, 26, 29, 51, 53, 57]. Before we continue, we should mention the main weakness of M-bases compared to Schauder bases: while it remains true that every vector $x\in\mathcal{X}$ can be unambiguously identified with its coordinates $(\langle\varphi_{k},x\rangle)_{k=1}^{\infty}$, the series $\sum_{k=1}^{\infty}\langle\varphi_{k},x\rangle e_{k}$ need not converge in general.
@@
However, it is the non-separable framework the one where M-bases have proved being the most fruitful. As a first instance of this, let us mention that every Banach space with an M-basis admits a continuous injection in $c_{0}(\Gamma)$ and therefore a rotund renorming, see, e.g., [28, Theorem II.2.4]. In order to achieve stronger renorming properties one frequently requires additional assumptions on the M-basis. Perhaps the most important example is the celebrated Troyanski’s renorming theorem [107] that every Banach space with a strong M-basis admits a locally uniformly rotund (LUR, for short) norm. Let us also refer to [50, Theorem 3.48] for a proof and to [24], [87], and [88, Lemma 2.1] for related results. Among all properties of M-bases the closest one to renorming theory is that of a norming M-basis, since the very definition of norming M-basis involves an equivalent norm induced by the M-basis itself (see Section 2 for details).
@@
A related use of M-bases in non-separable Banach spaces is to characterise several classes of non-separable Banach spaces by the existence of M-bases with certain additional properties. The most important examples of this are weakly Lindelöf determined (WLD, for short) and Plichko spaces that we will define in Section 3 and will play a role in several parts of the paper. Let us refer to [50, Chapter 6] and [115] for more on this perspective; let us also mention in passing a parallel line of research aiming at characterising classes of Banach spaces via certain properties of so called projectional skeletons, see, e.g., [20, 23, 39].
@@
One of the main open problems in the area, still widely open, has been to detect which class of Banach spaces is characterised by the existence of a norming M-basis. It has been conjectured in the Seventies that there ought to be a connection between norming M-bases and weakly compactly generated Banach spaces (WCG for short). This resulted in a famous open problem, due to John and Zizler [60], whether every WCG Banach space admits a norming M-basis and to a question in the opposite direction, due to Godefroy [46, Problem 112], whether Asplund spaces with norming M-bases are WCG. Both these problems have been recently solved in the negative [49, 52] and we shall give more details on this in Section 4. On the other hand, we also prove in Section 4 that the answer to Godefroy’s problem is positive under the stronger assumption that $\mathcal{X}$ has countable Szlenk index (Theorem 4.16).
@@
is easily seen to be a continuous injection (here, $[-1,1]^{\mathbb{N}}$ is endowed with the product topology). By the compactness of $(B_{\mathcal{X}^{*}},w^{*})$, it is a homeomorphic embedding, which implies the standard fact that $(B_{\mathcal{X}^{*}},w^{*})$ is metrisable. The same evaluation map $\mathsf{ev}$, now having values in $[-1,1]^{\Gamma}$, can be defined in presence of any M-basis and it still is a homeomorphic embedding. Depending on the range of $\mathsf{ev}$ one can classify several classes of Banach spaces: for example, when $\mathcal{X}$ is WLD $\mathsf{ev}(\varphi)$ only has countably many non-zero coordinates, which leads to the notion of a Corson compact space. Similar results are available when $\mathcal{X}$ is WCG, Hilbert generated, or Plichko, as we will explain in Section 3. When the M-basis is $1$-norming, the description of the image of $\mathsf{ev}$ leads to the notion of semi-Eberlein compact space, first introduced in [78]. In Section 7 we shall define this class of compact spaces, describe some known properties, and state several open problems.
@@ 2. Examples of M-bases
For the sake of completeness, we begin by explicitly stating the definition of an M-basis without the separability constraint. Given a Banach space $\mathcal{X}$, a family $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}\subseteq\mathcal{X}\times\mathcal{X}^{*}$ is a biorthogonal system in $\mathcal{X}$ if $\langle\varphi_{\alpha},e_{\beta}\rangle=\delta_{\alpha,\beta}$ for all $\alpha,\beta\in\Gamma$. The biorthogonal system $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is fundamental (or complete) if ${\rm span}\{e_{\alpha}\}_{\alpha\in\Gamma}$ is dense in $\mathcal{X}$ and it is total if ${\rm span}\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is $w^{*}$-dense in $\mathcal{X}^{*}$ (equivalently, if it separates points of $\mathcal{X}$: the unique vector $x\in\mathcal{X}$ with $\langle\varphi_{\alpha},x\rangle=0$ for all $\alpha\in\Gamma$ is $x=0$). A Markushevich basis (henceforth, M-basis) is a fundamental and total biorthogonal system. As we said in the Introduction, one frequently requires M-bases with additional properties, that are frequently obtained by strengthening the sense in which the functionals $\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ exhaust the dual. The strongest notion is that of shrinking M-basis, obtained by requiring that ${\rm span}\{\varphi_{\alpha}\}_{\alpha\in\Gamma}$ is dense in $\mathcal{X}^{*}$ (in the norm topology); as we will see in Theorem 4.3 this is a rather restrictive definition.
@@
Observe that $\mathcal{Z}$ becomes a $1$-norming subspace in the norm $\left\|\cdot\right\|_{\mathcal{Z}}$. Let us refer to [47, Section 3.4] for some additional elementary remarks on the notion of norming subspaces.
@@
$\int_{\mathbb{T}}\alpha F_{n}(t_{0}-t)f(t)\ {\rm d}\mu(t)=\alpha(F_{n}*f)(t_{0})\to\|f\|_{\infty}.$
@@
More generally, $\mathcal{C}(\mathcal{K})$ has a $1$-norming M-basis whenever $\mathcal{K}$ is a compact Abelian group, [73, Theorem 4.6].
@@
Every Schauder basis is a norming M-basis and every monotone Schauder basis is a $1$-norming one. This is elementary and we refer to [47, Remark 32(3)] for a proof. A similar argument shows that every long unconditional basis is a norming M-basis as well, [50, Exercise 7.8]. In particular, the canonical bases of $c_{0}(\Gamma)$ and $\ell_{p}(\Gamma)$, $1\leqslant p<\infty$ are $1$-norming. In the non-separable case the unconditionality assumption is essential, as not every long Schauder basis is norming (see Remark 5.5).
In the opposite direction, the classical Mazur technique shows that every norming M-basis admits a Schauder subsequence; we state this only for countable M-bases, but a small variation gives the result for uncountable ones as well, [52, Fact 4.4].
@@
and a standard inductive argument (as in [38, Theorem 4.19]) finishes the proof. ∎
Next, we discuss the existence of (norming) M-bases in separable Banach spaces. The following crucial result, essentially based on the Gram-Schmidt orthogonalisation algorithm, is due to Markushevich [85]. We refer to [38, Theorem 4.59] for a proof.
@@
${\rm span}\{e_{k}\}_{k=1}^{\infty}={\rm span}\{x_{k}\}_{k=1}^{\infty}\qquad\mbox{and}\qquad{\rm span}\{\varphi_{k}\}_{k=1}^{\infty}={\rm span}\{\psi_{k}\}_{k=1}^{\infty}.$
The result immediately implies the existence of a $1$-norming M-basis in every separable Banach space. Indeed, it is enough to take a dense sequence $(x_{k})_{k=1}^{\infty}$ in $\mathcal{X}$, let $\psi_{k}$ be a norm one functional with $\langle\psi_{k},x_{k}\rangle=1$ and observe that ${\rm span}\{\psi_{k}\}_{k=1}^{\infty}$ is $1$-norming for $\mathcal{X}$.
The situation is radically different for non-separable $\mathcal{X}$ and there are non-separable Banach spaces without an M-basis, the simplest example of this being $\ell_{\infty}$, [65]. It is even consistent with ZFC that there are non-separable Banach spaces without uncountable biorthogonal systems, [50, Section 4.4].
Part of the interest of Theorem 2.5 is in the clause concerning the preservation of the linear spans, that allows for instance to obtain examples of non-norming M-bases.
@@
When searching for M-bases that fail to be norming one immediately faces the problem of finding $w^{*}$-dense subspaces that are not norming. The existence of such subspaces has been clarified in [25]: A Banach space $\mathcal{X}$ has a $w^{*}$-dense not norming subspace in $\mathcal{X}^{*}$ if and only if $\mathcal{X}^{**}/\mathcal{X}$ is infinite-dimensional (i.e., $\mathcal{X}$ is not quasi-reflexive). Therefore, by Theorem 2.5 every separable Banach space that is not quasi-reflexive admits a non-norming M-basis. Conversely, in a separable quasi-reflexive Banach space all M-bases are norming (and in a separable reflexive space all M-bases are shrinking).
@@
For instance, the trigonometric system of Example 2.2 is strong. Despite how innocent this condition might look there are M-bases that fail to be strong and it is actually rather easy to find one in every separable Banach space, [50, Proposition 1.34]. Inspection of this proof shows that, if $\{e_{k};\varphi_{k}\}_{k=1}^{\infty}$ is an M-basis for $\mathcal{X}$, then there is another M-basis for $\mathcal{X}$ with the same linear spans and that fails to be strong. Starting from a norming M-basis shows that every separable Banach space admits a norming M-basis that is not strong. We will see an example of the opposite situation in Remark 5.6.
@@ 3. Classes of non-separable Banach spaces and M-bases
WCG Banach spaces admit a characterisation in terms of M-bases as follows: a Banach space $\mathcal{X}$ is WCG if and only if it admits a weakly compact M-basis, i.e., an M-basis $\{e_{\alpha};\varphi_{\alpha}\}_{\alpha\in\Gamma}$ such that $\{e_{\alpha}\}_{\alpha\in\Gamma}\cup\{0\}$ is weakly compact in $\mathcal{X}$ (see [50, Theorem 6.9]). It is easy to reformulate this property in terms of the evaluation map $\mathsf{ev}$: the M-basis is weakly compact if and only if
@@
When $\mathcal{X}$ is a $\mathcal{C}(\mathcal{K})$ space, the situation is more symmetric: in fact, $\mathcal{C}(\mathcal{K})$ is WCG if and only if $\mathcal{K}$ is Eberlein, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is Eberlein, [5]. This and the fact that $\mathcal{X}$ is a subspace of $\mathcal{C}(B_{\mathcal{X}^{*}})$ imply that, when $B_{\mathcal{X}^{*}}$ is Eberlein, $\mathcal{X}$ is a subspace of the WCG space $\mathcal{C}(B_{\mathcal{X}^{*}})$. Therefore the dual ball of $\mathcal{X}$ is Eberlein if and only if $\mathcal{X}$ is a subspace of a WCG space.
At one point, we will need a subclass of the class of WCG spaces, namely Hilbert generated Banach spaces and the corresponding class of uniform Eberlein compacta. A Banach space $\mathcal{X}$ is Hilbert generated if there are a Hilbert space $\mathcal{H}$ and a linear map $T\colon\mathcal{H}\to\mathcal{X}$ with dense range. By definition, separable Banach spaces, $c_{0}(\Gamma)$ and $L_{1}(\mu)$ for a finite measure $\mu$ are Hilbert generated. However, there are reflexive Banach spaces that are not Hilbert generated [79] and even super-reflexive ones, [35, Section 4]. Since Hilbert spaces are WCG, every Hilbert generated Banach space is WCG. Dually, a compact topological space $\mathcal{K}$ is uniform Eberlein if it is homeomorphic to a weakly compact subset of a Hilbert space. Analogous results as those of the previous paragraphs give that $\mathcal{X}$ is Hilbert generated if and only if
@@
and that $\mathcal{X}$ is a subspace of a Hilbert generated space if and only if $(B_{\mathcal{X}^{*}},w^{*})$ is uniform Eberlein. Moreover, a $\mathcal{C}(\mathcal{K})$ space is Hilbert generated if and only if $\mathcal{K}$ is uniform Eberlein, if and only if $(B_{\mathcal{C}(\mathcal{K})^{*}},w^{*})$ is uniform Eberlein, [12, Theorem 3.2]. For further details we refer to [50, Section 6.3] and [12, 35, 37].
@@
In this case, it turns out that every M-basis countably supports $\mathcal{X}^{*}$, [50, Theorem 5.37]. Comparing this notion to the M-basis characterisation of WCG spaces one easily sees that every WCG Banach space is WLD. Moreover, WLD spaces are closed under quotients and subspaces, as we will mention below (thus subspaces of WCG are also WLD). It is not easy to offer natural examples of WLD spaces that are not WCG; one could mention for instance Rosenthal’s non WCG subspace of a WCG space [94], or Talagrand’s example of a WLD space that is not subspace of a WCG space, [103].
@@
Therefore, $(B_{\mathcal{X}^{*}},w^{*})$ satisfies the following definition. A compact topological space $\mathcal{K}$ is Corson if it is homeomorphic to a subset of $\Sigma(\Gamma)$, for some set $\Gamma$. Since elements in $c_{0}(\Gamma)$ are countably supported, every Eberlein compact space is Corson. Moreover, Corson compacta are closed under subspaces and continuous images, [45, 48, 86]. As it turns out, a Banach space is WLD if and only if its dual ball is Corson, [7, Section 1], [111, Theorem 1.1] and references therein. Therefore, we get in particular that WLD spaces are closed under quotients and subspaces. Finally, a $\mathcal{C}(\mathcal{K})$ space is WLD if and only if $\mathcal{K}$ is Corson and it has property (M), namely every measure on $\mathcal{K}$ has separable support, [8].
At one point (in Theorem 4.3) we will need one more class, intermediate between WCG and WLD Banach spaces. In order not to overload the section with definitions we prefer not to enter its definition and only give some references. The class consists of the so-called weakly countably determined (WCD, for short), or Vašák, Banach spaces; as we said, we have the implications WCG $\implies$ WCD $\implies$ WLD. WCD spaces were introduced in [112], see e.g., [28, Section VI.2], or [47, p. 174 and p. 433].